# Inference of entropy production for periodically driven systems

### v1

@@ I Introduction
It is often the case in an experiment that not all states of the mesoscopic system are accessible. Therefore, the inference of properties of the system, such as its rate of entropy production, from partial information is a fundamental theoretical problem in stochastic thermodynamics. In particular, while well controlled experiments with full information about the system have provided beautiful connections between modern theory and experiment Collin et al. (2005); Blickle et al. (2009); Bérut et al. (2012); Koski et al. (2014); Dieterich et al. (2015); Kumari et al. (2020), connecting stochastic thermodynamics with less controlled experiments in molecular biophysics remains a major challenge, which is closely related to the problem of inferring properties of the system from partial information.
Much work on the problem of inferring the entropy production and related observables has been done. The thermodynamic uncertainty relation Barato and Seifert (2015a); Gingrich et al. (2016); Pietzonka et al. (2016a) has been used to infer entropy production and related observables from the fluctuations of a current Barato and Seifert (2015b); Pietzonka et al. (2016b, 2017); Li et al. (2019); Van Vu et al. (2020); Dechant and Sasa (2021). The rate of entropy production can also be estimated from the Kullback-Leibler divergence between the probability of a coarse-grained trajectory and the probability of the respective reversed trajectory Roldan and Parrondo (2010, 2012); Muy et al. (2013); Martínez et al. (2019). Another approach to estimate the entropy production is to assume that the visible dynamics corresponds to a hidden Markov process and use the observable data to infer the underlying Markov process Ehrich (2021); Skinner and Dunkel (2021a, b). More generally, the related topic of coarse-graining in stochastic thermodynamics has been widely studied Rahav and Jarzynski (2007); Puglisi et al. (2010); Esposito (2012); Mehl et al. (2012); Bo and Celani (2014); Knoch and Speck (2015); Shiraishi and Sagawa (2015); Bisker et al. (2017); Seiferth et al. (2020); Teza and Stella (2020). Most of these works on estimation of entropy production are suitable for steady states, however, some works have recently addressed time-dependent cases Degünther et al. (2024); Singh and Proesmans (2023); Lee et al. (2023); Otsubo et al. (2022).
Particularly important for this work is a method to estimate the rate of the entropy production from the statistics of the sequences and times between a few visible transitions proposed independently in two recent papers van der Meer et al. (2022); Harunari et al. (2022). This approach is promising as it is reasonable to expect that the statistics of the time between transitions, i.e., the intertransition time, is accessible in an experiment. These two references concentrate on systems in nonequilibrium steady states, which are driven by a constant thermodynamic force.
@@
In this paper we provide answers to these questions. We show that the rate of entropy production can be inferred from the statistics of sequences and intertransition times between a few visible transitions in periodically driven systems. We find several important differences in relation to steady states. First, for periodically driven systems, not only the intertransition time is relevant but the initial time of the first transition is also important given that transition rates are time-dependent. Here, we develop an estimator that only depends on the intertransition time, eliminating the need to monitor the protocol, which can be challenging in experimental setups. However, eliminating this time-dependence leads to an inequality between the estimator and the real rate of entropy production that contains an extra term. Second, the estimator of the entropy production involves two different distributions of the intertransition time, one associated with the original forward protocol and the other associated with the reversed protocol.
Concerning the application of the method, we find that it is necessary to generate trajectories with a time-reversed protocol, and its usefulness is more involved for periodically driven systems. We show this difference with a simple three-state model with time-dependent energies and energy barriers Ray and Barato (2017), which we use as a proof of concept. First, for a unicyclic network of states, while a similar method gives the exact value of the entropy production for steady states van der Meer et al. (2022); Harunari et al. (2022), our estimate only provides a lower bound on the rate of entropy production for periodically driven systems. Second, in contrast to steady states, periodically driven systems can have entropy production in the absence of a net current in the space of states. For this case of absence of net motion, the estimator for entropy production gives a much smaller number than the real value, in other words, the presence of a current seems to be an important ingredient.
An analytical procedure to calculate the distribution of the probability density of the intertransition time for steady states was introduced in van der Meer et al. (2022); Harunari et al. (2022). This procedure maps the problem of calculating the intertransition time distribution into the problem of the calculating the time-dependent probability of an absorbing state of a certain auxiliary process. We generalize this procedure to periodically driven system, providing a pathway to obtain our estimator analytically.
@@ II.2 Model definition
As a proof of concept we consider the following model for a molecular pump depicted in Fig. 1. A similar model, with a stochastic protocol instead of the deterministic protocol considered here has been analyzed in Ray and Barato (2017). The model has three states, which we label as $i=1,2,3$. There is one energy $F_{e}$ and one energy barrier $B$, while the remaining energies and energy barriers are zero. The periodic protocol is piece-wise, with the period divide into three parts. For the first part, corresponding to $30<t<T/$ 3, the energy of state $1$ is $E_{1}(t)=F_{e}$ and the energy barrier between states 3 and 1 is $B_{31}(t)=B$. The other two energies and two energy barriers are 00. For the second part, corresponding to $T/3<t<2T/3$, $E_{2}(t)=F_{e}$, $B_{12}(t)=B$, and the other energies and energy barriers are zero. For the third part, corresponding to $2T/3<t<T$, $E_{3}(t)=F_{e}$, $B_{23}(t)=B$, and the other energies and energy barriers are zero.
@@
In other words, if $B=0$ then $J=0$, however, the periodically driven system is still out of equilibrium for $F_{e}\neq 0$ with a non-zero entropy production $\sigma$. In stark contrast to steady states, for which non-zero entropy production implies non-zero currents. As shown below, the presence of this net motion quantified by $J$ in periodically driven systems is essential for our estimate of the rate the entropy production. If $J=0$ our estimate is much smaller than the entropy production in general, providing little to no information.
We also consider the case of a the presence of fixed thermodynamic affinity $F$ that would drive the system to a non-equilibrium steady state in the absence of a periodic protocol. The transition rates are modified to
@@ II.3 Intertransition time
An example of a stochastic trajectory is shown in Fig. 1. This trajectory is a sequence of transitions and waiting times between transitions. We denote a jump, or transition, from state $i$ to state $j$ as $\ell=ij$. The reversed jump from $j$ to $i$ is denoted by $\bar{\ell}$. The set of visible transitions is denoted by $\mathcal{L}$. For instance, for the three-state model we have a total of six possible transitions, two between each pair of sates. If we only have access to the two transitions between the pair of sates $1$ and $2$, then two of the six transitions are part of the set $\mathcal{L}$.
@@ III Results for case study
The numerical results shown in the figures in this Section were obtained via Monte Carlo simulation with discretized time, where the time-step, which is much shorter than the characteristic time of the fastest jump, is small enough such that no relevant difference is observed by taking a smaller time-step. The statistics of sequences of visible jumps and their intertransition times were obtained by analyzing their frequency from long enough trajectories. The parameter $k$, which sets the time-scale of the transition rates in Eq. (8) and Eq. (9), and the period $T$ are set to $k=10$ and $T=1$. We here consider that only the transitions between states $1$ and $2$ are visible, i.e., =$\mathcal{L}=\{1\to 2,2\to 1\}$.
@@ III.1 Energy, energy barrier and entropy production estimate
Figure 3: Entropy production $\sigma$ and its estimate $\hat{\sigma}$ as functions of (a) the energy barrier $F_{e}$ for $B=0$ and as a functions of (b) the energy $F_{e}$ for fixed energy barrier $B=0$.
@@ III.2 Role of external fixed affinity $F$
Figure 4: Entropy production $\sigma=FJ+F_{e}J_{e}$ and the estimate $\hat{\sigma}$ as functions of the energy energy $F_{e}$ for (a) $B=0$ and for (b) $B=3$.
@@
In Fig. 4 we show our results for $B\neq 0$. Here, we can see that the estimate $\hat{\sigma}$ is sensitive to the work term $J_{e}F_{e}$, even in the regime where $-JF$ becomes negative. In this regime, the systems operates as an engine with work exerted against the internal force $F$. Therefore, the results in Fig. 4, together with the need for net motion, suggest that the presence of energy barriers seem to be a necessary condition for the estimate $\hat{\sigma}$ to capture information on the work term $F_{e}J_{e}$ of the rate of entropy production $\sigma$.
III.3 Extra term $X$
We did not find a direct inequality connecting the estimate $\hat{\sigma}$ and the rate of entropy production $\sigma$. However, in all our numerical results we observed the estimate $\hat{\sigma}$ below $\sigma$. A formal inequality $\sigma\geq\hat{\sigma}-X$ can be obtained with an extra term $X\geq 0$, as demonstrated in the next section. In Fig. 3 and Fig. 4 we also illustrate this inequality by plotting $\hat{\sigma}-X$, which can be negative.
@@ IV Proof of the bound and the emergence of the extra term
Notice that a proper time-reversal operation has to commute with coarse-graining Hartich and Godec (2021); Bisker et al. (2022), thus we consider maps satisfying $(\theta\gamma)^{\dagger}=\theta(\gamma^{\dagger})\eqqcolon\Gamma^{\dagger}$. A lower bound on entropy production rate $\sigma$, accessible from the statistics of the visible transition that pertain to $\mathcal{L}$, is defined as
@@
$P^{\dagger}[\Gamma^{\dagger}]=\mathcal{P}^{\dagger}_{\ell_{N},t_{f}-t_{N}}\prod_{i=0}^{N-2}\mathcal{P}^{\dagger}_{\overline{\ell}_{N-i-1},t_{f}-t_{N-i-1}|\overline{\ell}_{N-i},t_{f}-t_{N-i}}.$ (22)
@@
where $\mathcal{P}_{\ell^{\prime},t^{\prime};\ell,t}=\mathcal{P}_{\ell^{\prime},t^{\prime}|\ell,t}\mathcal{P}_{\ell,t}$ is the joint distribution. In this expression, since we are considering the limit $t_{f}\to\infty$, the contribution of the boundary term in $\ln(P[\Gamma_{t_{f}}]/P^{\dagger}[\Gamma_{t_{f}}^{\dagger}])$, which is $\ln(\mathcal{P}_{t_{1},\ell_{1}}/\mathcal{P}^{\dagger}_{t_{f}-t_{N},\ell_{N}})$, goes to zero.
@@
$\sigma_{\mathcal{L}}=\frac{K}{T}\sum_{\ell^{\prime},\ell\in\mathcal{L}}\int_{0}^{\infty}\mathrm{d}\tau\int_{0}^{T}\mathrm{d}t_{0}\mathcal{P}_{\ell^{\prime},{\tau+t_{0}};\ell,t_{0}}\ln\frac{\mathcal{P}_{\ell^{\prime},{\tau+t_{0}}|\ell,t_{0}}}{\mathcal{P}^{\dagger}_{\overline{\ell},\tau+t_{0}^{*}|\overline{\ell},t_{0}^{*}}},$ (24)
where $t_{0}^{*}=T-[(\tau+t_{0})\mod T]$. This $\sigma_{\mathcal{L}}$ can provide a better estimate of $\sigma$ in comparison $\hat{\sigma}$. However, in practice it would require histograms that depends on two times, the initial time of the first transition and the intertransition time, which compromises its applicability.
@@
$X\equiv\frac{K}{T}\sum_{\ell^{\prime},\ell\in\mathcal{L}}\int_{0}^{\infty}\mathrm{d}t\int_{0}^{T}\mathrm{d}t_{0}P_{\ell^{\prime},{\tau+t_{0}};\ell,t_{0}}\ln\frac{P_{t_{0}|\ell}}{P^{\dagger}_{t_{0}^{*}|\overline{\ell^{\prime}}}}.$ (25)
From this definition of $X$ and (24), using the log-sum inequality in $t_{0}$ we obtain $\sigma_{\mathcal{L}}+X\geq\hat{\sigma}$, where $\hat{\sigma}$ is our estimate given by Eq. (14). Since $\sigma\geq\sigma_{\mathcal{L}}$ we obtain
@@
The extra terms satisfies $X\geq 0$, this inequality can be obtained from Eqs. (12), (25), and the log-sum inequality in $t_{0}$. The extra term can be calculated from a trajectory (and another one with the reversed protocol) in the following way. First we obtain the distributions $\mathcal{P}_{t_{0}|\ell}$ from a trajectory with forward protocol and $\mathcal{P}^{\dagger}_{t_{0}|\ell}$ from a trajectory with reversed protocol. Then we can run over all transitions in a visible trajectory $\Gamma$ with forward protocol, with its probability represented in Eq. (21), and compute $\ln(\mathcal{P}_{t_{i}|\ell_{i}}/P^{\dagger}_{t_{f}-t_{i}|\overline{\ell}_{i}})$ whenever a visible transition $\ell_{i}$ happens at time $t_{i}$. If the probabilities are empirically inferred from the same trajectory, correlations can give rise to convergence issues, thus it is better to use independent trajectories. Note that there is no need to calculate a distribution that depends on two times in order to calculate $X$ from a trajectory.
In summary, our estimate $\hat{\sigma}$ is not connected to $\sigma$ by a formal inequality to our knowledge. However, within our numerics $\hat{\sigma}$ is below $\sigma$ in all cases. If we consider $\hat{\sigma}-X$ as an estimate, then we do have a formal inequality.

### v2

@@ I Introduction
It is often the case in an experiment that not all states of the mesoscopic system are accessible. Therefore, the inference of properties of the system, such as its rate of entropy production, from partial information is a fundamental theoretical problem in stochastic thermodynamics. In particular, while well-controlled experiments with full information about the system have provided beautiful connections between modern theory and experiment Collin et al. (2005); Blickle et al. (2009); Bérut et al. (2012); Koski et al. (2014); Dieterich et al. (2015); Kumari et al. (2020), connecting stochastic thermodynamics with less controlled experiments in molecular biophysics remains a major challenge, which is closely related to the problem of inferring properties of the system from partial information.
Much work on the problem of inferring the entropy production and related observables has been done. The thermodynamic uncertainty relation Barato and Seifert (2015a); Gingrich et al. (2016); Pietzonka et al. (2016a) has been used to infer entropy production and related observables from the fluctuations of a current Barato and Seifert (2015b); Pietzonka et al. (2016b, 2017); Li et al. (2019); Van Vu et al. (2020); Dechant and Sasa (2021). The rate of entropy production can also be estimated from the Kullback-Leibler divergence between the probability of a coarse-grained trajectory and the probability of the respective reversed trajectory Roldan and Parrondo (2010, 2012); Muy et al. (2013); Martínez et al. (2019). Another approach to estimate the entropy production is to assume that the visible dynamics corresponds to a hidden Markov process and use the observable data to infer the underlying Markov process Ehrich (2021); Skinner and Dunkel (2021a, b). More generally, the related topic of coarse-graining in stochastic thermodynamics has been widely studied Rahav and Jarzynski (2007); Puglisi et al. (2010); Esposito (2012); Mehl et al. (2012); Bo and Celani (2014); Knoch and Speck (2015); Shiraishi and Sagawa (2015); Bisker et al. (2017); Seiferth et al. (2020); Teza and Stella (2020). Most of these works on estimation of entropy production are suitable for steady states, however, some works have recently addressed time-dependent cases Degünther et al. (2024); Singh and Proesmans (2023); Lee et al. (2023); Otsubo et al. (2022); Koyuk et al. (2020).
Particularly important for this work is a method to estimate the rate of the entropy production from the statistics of the sequences and times between a few visible transitions proposed independently in two recent papers van der Meer et al. (2022); Harunari et al. (2022). This approach is promising as it is reasonable to expect that the statistics of the time between transitions, i.e., the intertransition time, is accessible in an experiment. These two references concentrate on systems in nonequilibrium steady states, which are driven by constant thermodynamic forces.
@@
In this paper we provide answers to these questions. We show that the rate of entropy production can be inferred from the statistics of sequences and intertransition times between a few visible transitions in periodically driven systems. We find several important differences in relation to steady states. Since transition rates are time-dependent, the statistics of transition pairs depend on the intertransition time and the initial time of the first transition. We overcome this issue by developing an estimator that only depends on the intertransition time. The formal elimination of this dependence on the time of the first transition leads to an inequality between the estimator and the real rate of entropy production that contains an extra term. We also find an estimator that depends on both the intertransition time and the time of the first transition. Even though this bound provides a better estimate of the rate of entropy production, compared to the one that only depends on intertransition times, in practice, it requires one histogram of intertransition times per each considered initial time, which can be computationally expensive.
The estimator of the entropy production involves two different distributions of the intertransition time, one associated with the original forward protocol and the other associated with the time-reversed protocol. In contrast to steady states that only involve one distribution. Hence, the application of the method requires two “experiments” one with forward protocol and another with the time-reversed protocol.
A simple three-state model with time-dependent energies and energy barriers Ray and Barato (2017) is used as a proof of concept. We obtain the following results with this model. First, for a unicyclic network of states, while a similar method gives the exact value of the entropy production for steady states van der Meer et al. (2022); Harunari et al. (2022), our estimate only provides a lower bound on the rate of entropy production for periodically driven systems. Second, in contrast to steady states, periodically driven systems can have entropy production in the absence of a net current in the space of states. For this case of absence of net motion, the estimator for entropy production gives a numerical result compatible with zero, which suggests that net motion is an essential feature for our estimator to provide information about the rate of entropy production.
An analytical procedure to calculate the distribution of the probability density of the intertransition time for steady states was introduced in van der Meer et al. (2022); Harunari et al. (2022); Sekimoto (2021). This procedure maps the problem of calculating the intertransition time distribution into the problem of calculating the time-dependent probability of an absorbing state of a certain auxiliary process. We generalize this procedure to periodically driven system, providing a pathway to obtain our estimator analytically.
@@ II.2 Model definition
As a proof of concept we consider the following model for a molecular pump depicted in Fig. 1. A similar model, with a stochastic protocol instead of the deterministic protocol considered here has been analyzed in Ray and Barato (2017). The model has three states, which we label as $i=1,2,3$. There is one energy $F_{e}$ and one energy barrier $B$, while the remaining energies and energy barriers are zero. The periodic protocol is piece-wise, with the period divided into three parts. For the first part, corresponding to $30<t<T/$ 3, the energy of state $1$ is $E_{1}(t)=F_{e}$ and the energy barrier between states 3 and 1 is $B_{31}(t)=B$. The other two energies and two energy barriers are 00. For the second part, corresponding to $T/3<t<2T/3$, $E_{2}(t)=F_{e}$, $B_{12}(t)=B$, and the other energies and energy barriers are zero. For the third part, corresponding to $2T/3<t<T$, $E_{3}(t)=F_{e}$, $B_{23}(t)=B$, and the other energies and energy barriers are zero.
@@
In other words, when $B=0$, $J$ also vanishes. However, even with $J=0$, the periodically driven system remains out of equilibrium when $Fe\neq 0$, resulting in a positive entropy production. This behavior contrasts sharply with steady state systems, where nonzero entropy production is always associated with currents. As shown below, the presence of this net motion quantified by $J$ in periodically driven systems is essential for our estimate of the rate the entropy production. If $J=0$ our estimate is much smaller than the entropy production in general, providing little to no information.
We also consider the case of the presence of a fixed thermodynamic affinity $F$ that would drive the system to a non-equilibrium steady state in the absence of a periodic protocol. The transition rates are modified to
@@ II.3 Intertransition time
An example of a stochastic trajectory is shown in Fig. 1. This trajectory is a sequence of transitions and waiting times between transitions. We denote a jump, or transition, from state $i$ to state $j$ as $\ell=ij$. The reversed jump from $j$ to $i$ is denoted by $\bar{\ell}$. The set of visible transitions is denoted by $\mathcal{L}$. For instance, for the three-state model we have a total of six possible transitions, two between each pair of states. If we only have access to the two transitions between the pair of states $1$ and $2$, then two of the six transitions are part of the set $\mathcal{L}$.
@@ III Results for case study
In this section, we show our numerical study of entropy production and the estimator for the model defined in Sec. II.2. We observed the estimate $\hat{\sigma}$ below the real rate of entropy production $\sigma$ in all our numerical results within this model. However, the formal inequality connecting both quantities contains an extra term $X$ and reads $\sigma\geq\hat{\sigma}-X$, where the extra term fulfills $X\geq 0$. This inequality is proved in the next section. In Fig. 3 and Fig. 4 we also illustrate it by plotting $\hat{\sigma}-X$, which can be negative.
The numerical results shown in the figures in this Section were obtained via Monte Carlo simulation with discretized time, where the time-step, which is much shorter than the characteristic time of the fastest jump, is small enough such that no relevant difference is observed by taking a smaller time-step. The statistics of sequences of visible jumps and their intertransition times were obtained by analyzing their frequency from long enough trajectories such that the empirically measured quantities presented reasonable convergence. The parameter $k$, which sets the time-scale of the transition rates in Eq. (8) and Eq. (9), and the period $T$ are set to $k=10$ and $T=1$. We here consider that only the transitions between states $1$ and $2$ are visible, i.e., =$\mathcal{L}=\{1\to 2,2\to 1\}$.
@@ III.1 Energy, energy barrier and entropy production estimate
Figure 3: Entropy production $\sigma$, its estimate $\hat{\sigma}$ and its lower bound $\hat{\sigma}-X$ as functions of (a) the energy $F_{e}$ for $B=0$ and (b) the energy barrier $B$ for $F_{e}=2$.
@@ III.2 Role of external fixed affinity $F$
Figure 4: Entropy production $\sigma=FJ+F_{e}J_{e}$, its estimate $\hat{\sigma}$ and its lower bound $\hat{\sigma}-X$ as functions of (a) the energy $F_{e}$ for $B=0$ and (b) the energy barrier $B$ for $F_{e}=3$.
@@
In Fig. 4 we show our results for $B\neq 0$. Here, we can see that the estimate $\hat{\sigma}$ is sensitive to the work term $J_{e}F_{e}$, even in the regime where $-JF$ becomes negative. In this regime, the systems operates as an engine with work exerted against the internal force $F$. Therefore, the results in Fig. 4, together with the need for net motion, suggest that the presence of energy barriers seems to be a necessary condition for the estimate $\hat{\sigma}$ to capture information on the work term $F_{e}J_{e}$ of the rate of entropy production $\sigma$.
@@ IV.1 First bound $\sigma_{\mathcal{L}}$
We consider a time-reversal operation that commutes with the coarse-graining map, $(\theta\gamma)^{\dagger}=\theta(\gamma^{\dagger})\eqqcolon\Gamma^{\dagger}$, which is achieved by noticing that transitions $\ell$ are replaced by $\bar{\ell}$ in the time-reversed trajectory. A lower bound on entropy production rate $\sigma$, accessible from the statistics of the visible transition that pertain to $\mathcal{L}$, is defined as
@@
$P^{\dagger}[\Gamma^{\dagger}]=\mathcal{P}^{\dagger}_{\bar{\ell}_{N},t_{f}-t_{N}}\prod_{i=0}^{N-2}\mathcal{P}^{\dagger}_{\overline{\ell}_{N-i-1},t_{f}-t_{N-i-1}|\overline{\ell}_{N-i},t_{f}-t_{N-i}}.$ (22)
@@
where $\mathcal{P}_{\ell^{\prime},t^{\prime};\ell,t}=\mathcal{P}_{\ell^{\prime},t^{\prime}|\ell,t}\mathcal{P}_{\ell,t}$ is the joint distribution. In this expression, since we are considering the limit $t_{f}\to\infty$, the contribution of the boundary term in $\ln(P[\Gamma_{t_{f}}]/P^{\dagger}[\Gamma_{t_{f}}^{\dagger}])$, which is $\ln(\mathcal{P}_{\ell_{1},t_{1}}/\mathcal{P}^{\dagger}_{\bar{\ell}_{N},t_{f}-t_{N}})$, goes to zero.
Figure 5: The quantities of Fig. 4 and $\sigma_{\mathcal{L}}$.
@@
$\sigma_{\mathcal{L}}=\frac{K}{T}\sum_{\ell^{\prime},\ell\in\mathcal{L}}\int_{0}^{\infty}\mathrm{d}\tau\int_{0}^{T}\mathrm{d}t_{0}\mathcal{P}_{\ell^{\prime},{\tau+t_{0}};\ell,t_{0}}\ln\frac{\mathcal{P}_{\ell^{\prime},{\tau+t_{0}}|\ell,t_{0}}}{\mathcal{P}^{\dagger}_{\overline{\ell},\tau+t_{0}^{*}|\overline{\ell^{\prime}},t_{0}^{*}}},$ (24)
where $t_{0}^{*}=T-[(\tau+t_{0})\mod T]$. In Fig. 5 we show that $\sigma_{\mathcal{L}}$ can provide a better estimate of $\sigma$ in comparison to $\hat{\sigma}$, where $\sigma_{\mathcal{L}}$ was calculated using the analytical methods of Section V. However, in practice, it is much more difficult to obtain this estimator from a trajectory generated in a simulation or experiment. The main difference is that, for each pair of transitions, this quantity requires one histogram for each initial time $t_{0}$ in the period, while $\hat{\sigma}$ uses only one histogram.
IV.2 Elimination of $t_{0}$
@@
$X\equiv\frac{K}{T}\sum_{\ell^{\prime},\ell\in\mathcal{L}}\int_{0}^{\infty}\mathrm{d}\tau\int_{0}^{T}\mathrm{d}t_{0}P_{\ell^{\prime},{\tau+t_{0}};\ell,t_{0}}\ln\frac{P_{t_{0}|\ell}}{P^{\dagger}_{t_{0}^{*}|\overline{\ell^{\prime}}}}\geq 0.$ (25)
From this definition of $X$ and (24), using the log-sum inequality in $t_{0}$ we obtain
$\sigma_{\mathcal{L}}+X\geq\hat{\sigma},$ (26)
where $\hat{\sigma}$ is our estimate given by Eq. (14). Since $\sigma\geq\sigma_{\mathcal{L}}$ we obtain
@@
The extra term $X$ fulfills the inequality
$X\geq 0,$ (28)
which is obtained from Eqs. (12), (25), and the log-sum inequality in $t_{0}$. The term $X$ can be calculated from a trajectory (and another one with the reversed protocol) in the following way. First, we obtain the distributions $\mathcal{P}_{t_{0}|\ell}$ from a trajectory with forward protocol and $\mathcal{P}^{\dagger}_{t_{0}|\ell}$ from a trajectory with reversed protocol. Then we can run over all transitions in a visible trajectory $\Gamma$ with forward protocol, with its probability represented in Eq. (21), and compute $\ln(\mathcal{P}_{t_{i}|\ell_{i}}/P^{\dagger}_{t_{f}-t_{i}|\overline{\ell}_{i}})$ whenever a visible transition $\ell_{i}$ happens at time $t_{i}$. If the probabilities are empirically inferred from the same trajectory, correlations can give rise to convergence issues, thus it is better to use independent trajectories. Note that there is no need to calculate a distribution that depends on two times in order to calculate $X$ from a trajectory, which was a main practical issue with the tighter estimate $\sigma_{\mathcal{L}}$.
In summary, our estimate $\hat{\sigma}$ is not connected to $\sigma$ by a formal inequality to our knowledge. However, within our numerics restricted to one case study $\hat{\sigma}$ is below $\sigma$. This evidence is not enough to make a conjecture but it leaves the possibility of an inequality open. If we consider $\hat{\sigma}-X$ as an estimate, then we do have a formal inequality.